OpenStudy (anonymous):

(28 - 7b)/ (b -4) * (1/ ( b + 10)

OpenStudy (anonymous):

Are you able to factorise 28-7b [which is the numerator of the first term]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What did you get when you factorised it?

OpenStudy (anonymous):

i don't know how to factorize the problem i kind of tried and got confused what did you get

OpenStudy (luigi0210):

Come on eb you got this!

OpenStudy (anonymous):

You told me that you knew how to factorise: 28-7b. I want to see what you got for just doing that?

OpenStudy (anonymous):

\[\frac{ 27-7b}{ b ^{2-}40 }\]

OpenStudy (anonymous):

You're getting all the knowledge that you've learnt mixed up. You're confusing yourself even more, when I asked you to factorise 28-7b. Could you do that step for me please?

OpenStudy (anonymous):

i forgot how to do that i thought you was talking about subtracting them

OpenStudy (anonymous):

If you factorise, you take a common factor for each of the terms. For EXAMPLE: Factorise 4+8x. The factor for the two terms is 4. So we take the 4 out and we get: \[4(1+2x)\] WHen you distibute or expand, you would get the same two terms as before.

OpenStudy (anonymous):

One more note- If you take out the common factor, you have would be changing the terms inside the brackets. If you factorise 4+8x and take out 4, you don't go like this: \[4(4+8x)\] That would give you a completely different result. You must account for the factor taken out and you must make it possible to fit everything together. What would be the factor for 28-7b then?

OpenStudy (anonymous):

you would be*

OpenStudy (anonymous):

7(4-1b)

OpenStudy (anonymous):

Well done. you don't need the 1 because 1b=b. It's the same thing. Okay if we also factorise the minus sign out, what would we get then?

OpenStudy (anonymous):

So for EXAMPLE, if we take out the minus sign in this example: 4(1-2x)=-4(2x-1), that would be the end result.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Could you show me what you get now when you take out the minus from: 7(4-b)?

OpenStudy (anonymous):

7(4b)

OpenStudy (anonymous):

Did you read my example properly?

OpenStudy (anonymous):

Factorise the minus also means take the minus out.

OpenStudy (anonymous):

I don't mean literally take it out. I mean factorise the minus sign out.

OpenStudy (anonymous):

Don't you see that if you do that, it wouldn't make sense mathematically if you literally took the minus sign out?

OpenStudy (anonymous):

-7(b-4)

OpenStudy (anonymous):

@eb3001 YES! Now you have this: \[\frac{-7(b-4)}{b -4} \times \frac{1}{b + 10}\]

OpenStudy (anonymous):

Now you can cancel the b-4 with the denominator and numerator.

OpenStudy (luigi0210):

@eb3001 Come on you're almost done! Just follow along with @Azteck

OpenStudy (anonymous):

alright \[-7*\frac{ 1 }{ b+10 }\] so you get this

OpenStudy (anonymous):

\[-7* \frac{ -7 }{ 7b+70 }\]

OpenStudy (anonymous):

\[\frac{ 7b }{ 7b+70}\]

OpenStudy (luigi0210):

Oops, you can't multiply the whole thing by -7, just the numerator. If it was -7/-7 then in that cause you could.

OpenStudy (anonymous):

\[\frac{ -7b }{ b+10 }\]

OpenStudy (luigi0210):

Where did that b come from?

OpenStudy (anonymous):

Um you were correct when you came to this stage: \[-7\times \frac{1}{b+10}\]

OpenStudy (anonymous):

The last step would just be putting the -7 on top of 1/b+10

OpenStudy (anonymous):

my bad \[\frac{ -7 }{ b+10 }\]

OpenStudy (anonymous):

Yep. Correct! Good job.

OpenStudy (luigi0210):

Ha, there you go eb :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No worries.